OpenStudy (anonymous):

Solve the linear system by graphing. Check your solution. 13. y= -x +4 y = 2x - 8 please help I would appreciate it^.^ Explain every detail of how to do this so Ican do this for other equations.

OpenStudy (anonymous):

\[y=-x+4[1]\] \[y=2x-8[2]\] Sub in the RHS of equation [1] into y. If you don't understand how you can do that, well in the first equation, the equation states that y equals -x+4. In other words, y can be written as "-x+4" (without the quotation marks) So the second equation will become this: \[-x+4=2x-8\] Collect like terms. That means gather the terms that look "alike". \[4+8=2x+x\] \[3x=12\] \[x=4\] You can find y by substituting that x-value into ONE of the two equations you were given in the question.

OpenStudy (anonymous):

Can you just explain the whole thing pz

OpenStudy (anonymous):

Oh the question says grraphing. Graph the lines and show me your diagram.

OpenStudy (anonymous):

graphing*

OpenStudy (anonymous):

So, draw the two lines above on one number plane.

OpenStudy (anonymous):

:p

OpenStudy (anonymous):

Just show me how to do it and I will know how to do it for the rest of it. Please because I need to sleep and its 1:15 AM

OpenStudy (anonymous):

Give me each step and explaination

OpenStudy (anonymous):

THat's your very first step....sketch the two lines above.

OpenStudy (anonymous):

I'm not even skipping any steps. I'm telling you to sketch it which you should be able to if you're attempting these types of questions.